6.2 g of K_3PO_4 are dissolved in 200 mL of soluti
Answer #1

formula for the normality is

N = (weight of solute / equivelent of solute ) x 1000 / volume of the solven tin mL

weigh tof solute = 6.2 g

equivelent weight = molecular weight / charge that can form = 212.1663 / 3 = 70.7221

volume of the solvent = 200 mL

plug in all these values in above equation

N = (6.2 / 70.7221) x 1000 / 200 mL

= 0.0876 x 5

= 0.438 N